1.Which of the following statements is true according to Keynesian monetary theory?

Group of answer choices

Stable business expectations means that a decrease in interest rates will increase planned investment.

The liquidity trap means that an increase in the money supply may not cause interest rates to fall.

Monetary policy is more effective in stimulating AD than fiscal policy.

The proper target for setting monetary policy is the supply of money.

2.

The money supply has a value of $12 trillion, the potential level of GDP is $18 trillion, and the velocity of money is 3. According to the quantity theory of money an increase in money supply of $6 trillion will ___ the price level from ___.

Group of answer choices

increase, 2 to 3

increase, 12 to 18

decrease, 2 to 1

decrease, 18 to 15.7
Answer #1

1) Option A is correct.

Option B is correct.

Option C is incorrect as fiscal policy is more effective than monetary policy to raise AD.

The target of monetary policy is the supply of money, interest rate and supply of bank credit. Thus we cannot say that it is only supply of money which makes this option incorrect.

2) Quantity theory of money says that Money supply * Velocity of money = Price * GDP

M * V = P * T

12 * 3 = P * 18

P = 2

When money supply rises by 6 trillion,

18 * 3 = P * 18

P = 3

Thus option A is correct.